OpenStudy (anonymous):

HOW WOULD U SIMPLIFY 2X OVER -8 i need to have it in a fraction

OpenStudy (jdoe0001):

2X OVER -8, sounds like a fraction already :|

OpenStudy (anonymous):

ik but i need it simplified

OpenStudy (hexagon001):

x/-4

OpenStudy (hexagon001):

the 2/-8 simplifies to 1/-4

OpenStudy (hexagon001):

are you clear skittles?

OpenStudy (hexagon001):

i use to love skittle sweets when i was a kid?

OpenStudy (hexagon001):

^random

OpenStudy (anonymous):

alright thank u cuz i have to find the Y intercept for -8y-2x=-48

OpenStudy (anonymous):

and that was random lol

jimthompson5910 (jim_thompson5910):

y-intercept -8y-2x=-48 -8y-2(0)=-48 ... plug in x = 0 -8y - 0 = -48 -8y = -48 y = -48/(-8) y = 6 So the y-intercept is (0, 6)

OpenStudy (anonymous):

awesome thank u so much :)

jimthompson5910 (jim_thompson5910):

np